Generalization of some theorems of Steinhaus in locally compact groups S. Basu, Calcutta, India

Keywords: Demi-sphere, density point, σ-finite outer measure, regularity of measure AMS Subject classification: 28A05 Abstract: The purpose of this paper is to generalize Theorems X and XI of Steinhaus [5] in locally compact Hausdorff topological groups in the light of transformations more general than those leading to the former.